Alanna0310:

The opening balance of one of the 31-day billing cycles for Clay's credit card was $3300, but after 15 days Clay made a payment of $1900 to decrease his balance, and it stayed the same for the remainder of the billing cycle. If his credit card's APR is 28%, how much more in interest would he pay for the billing cycle with the previous balance method than with the adjusted balance method?

Alanna0310:

Help Please!!

tigerlover:

Hello and Welcome to QuestionCove! My names tigerlover and I am here to help with what I can. ^^

tigerlover:

Given: 31 days = 7,400 15 days = payment of 4,900 16 days = 7,400 - 4,900 = 2,500 Method 1: 7,400 * 22% * 31/365 = 138.27 Method 2: adjusted balance 7,400 * 22% * 15/365 = 66.90 2,500 * 22% * 16/365 = 24.11 66.90 + 24.11 = 90.21 138.27 - 90.21 = 48.06 She would pay 48.06 more with the previous balance method than with the adjusted balance method. How about giving this a try? ^u^

Alanna0310:

Hey!! and ok cool thanks!!